OpenStudy (anonymous):

The equation (x + 5) + (y + 3) = 169 models the position and range of the source of a radio signal. Describe the position of the source and the range of the signals. ???

OpenStudy (anonymous):

\[(x + 5)^2 + (y + 3)^2 = 169 \] maybe ?

OpenStudy (anonymous):

i am sure is (x+5)^2 + ( y+3)^2

OpenStudy (anonymous):

because If the equation (x+5)2 +(y+3)2=169 models the position and range of the source of a radio signal the position of the source is (-5, -3) and the range of the signals is 13 units

OpenStudy (anonymous):

how can the position of the source be (-5,-3), how did you get your findings? I was going to say x is the position with a range of + or - 5 and y is the range of 3,-3)

OpenStudy (anonymous):

you have right i have notice it can be both ^^ . The equation (x + 5)2 + (y + 3)2 = 169 is the equation for a circle which in general form is written as (x + a)2 + (y + b)2 = r2, where a is the offset of the circle's center in the x-direction, b is the offset of the circle's center in the y-direction, and r is the circle's radius. Therefore, the center of the circle is (5,3), and the radius is √(169) = 13.

OpenStudy (anonymous):

thanks I understand

OpenStudy (anonymous):

medal ^^